# Current affiliation:]Université Paris-Saclay, CNRS, Laboratoire de Physique des Solides, 91405 Orsay, France

### v1

@@ III Results and discussions
The most prominent difference from the standard DFT+U results (presented in Fig. 6) is the emergence of a shallow electron-like band with its minimum located at the $\overline{\mathrm{K}}$} point, clearly visible on the right side in Fig. 2(b). In the corresponding Fermi surface map (Fig. 2(c)), this band forms a triangular contour around the $\overline{\mathrm{K}}$} point, closely reproducing the experimentally observed $\beta$-feature. The orbital character analysis indicates that this state has a substantial Co contribution. Its appearance only within the CPT framework, which treats the Hubbard $U$ beyond the mean-field level, provides strong evidence that the $\beta$-feature originates from strong electronic correlations, in line with our earlier conjecture. In contrast, the remaining bands in the CPT calculation are very similar to those from DFT+U, which already provides a satisfactory description of the corresponding experimental spectra. An exception is the spectral weight around 0.8 eV binding energy, which originates from the highest S-derived band (see Appendix B) and is not captured by our reduced CPT model.
@@
Fig. 3 (a) shows the Fermi surface map of Co1/3TaS2 along the $k_{z}$ direction, obtained by varying photon energy from 30 to 90 eV in a few-eV steps. The data were processed using the same procedure as described in Refs.  [10, 23]. The $k_{z}$ dispersion plot reveals a two-dimensional (2D-like) Fermi surface, $\alpha_{2}$, visible at $k_{x}\approx\pm 0.3\,\text{\AA }^{-1}$, which does not disperse with $k_{z}$ and shows only variations in intensity. A faint additional 2D-like Fermi surface, $\alpha_{1}$, can also be observed near $k_{x}\approx\pm 0.15\,\text{\AA }^{-1}$, although it is too weak to be traced across all $k_{z}$ values. Both features arise from the Ta $d$ bands of 2H-TaS2: the outer band ($\alpha_{2}$) corresponds to the antibonding state, while the inner ($\alpha_{1}$), weaker one corresponds to the bonding state. DFT modeling predicts that the bonding band should exhibit noticeable $k_{z}$ dispersion and cross the Fermi level between the $\Gamma$ and A points. This likely explains why the $\alpha_{1}$ band is not observed across all $k_{z}$ values. In Fig. 3(a), it is only visible when it lies above the Fermi level which corresponds to A point. Indeed, the ARPES intensity plots measured with photon energies between $h\nu=38$-5050 eV (see Fig. 6) show the bonding band ($\alpha_{1}$) below the Fermi level. This photon energy range corresponds to $k_{z}$ values below 4 $\AA ^{-1}$ in Fig. 3(a), which accounts for the disappearance of the $\alpha_{1}$ for low $k_{z}$ values. This $k_{z}$ dispersion agrees with the previously observed enhancement of $c$-axis electronic transport in TMDs upon Co intercalation, as discussed for Co1/3NbS2 in Refs. [10, 36]. The $\beta$-feature appears in Fig. 3(a) at both sides of the $\overline{\mathrm{K}}$ point, with its intensity increasing at higher photon energies, consistent with the behavior seen in Co1/3NbS2.
Fig. 3(b) shows the same Fermi surface as in panel (a), but calculated using the CPT method. The $\beta$-feature is again visible on both sides of the K point. Ta $d$ bands are also visible with antibonding band showing virtually no dispersion while bonding band exhibits strong $k_{z}$ dependence. We also notice the characteristic periodicity of two Brillouin zones resulting from the fact that there are two Co and two Ta layers per one unit cell. The similar effect was already reported by Weber et al. [52] in NbSe2.
Notably, we observed that the amount of Co intercalation strongly influences the presence of the $\beta$-feature. Fig. 4(a) shows the ARPES intensity plot of Co0.22TaS2. The measured band structure and Fermi surface resemble almost a rigid-band-shifted 2H-TaS2 band structure without the $\beta$-feature. By changing incident photon energy from 57 to 90 eV, in Fig. 4(b) we reconstructed the Fermi surface map along the $k_{z}$-direction. It shows a quasi-2D character with a small modulation, though no $\beta$-feature was observed within the experimental photon energy range. The central Fermi surface $\alpha$, around the $\mathrm{\Gamma}$ point, is observed at $k_{x}$ $\sim$ 0.5 Å-1 very close to the one in 2H-TaS2 [32] indicating a much smaller band shift compared to the 33% intercalated compound and suggesting even lower charge transfer. The same indication comes from the magnetic moment of Co1/3TaS2 compared to Co0.22TaS2. Co1/3TaS2 exhibits an effective magnetic moment of $\sim 3.3\,\mu_{B}$ similar to Co1/3NbS2 [36]. Despite being reduced compared to the spin-only value $\mu_{\mathrm{eff}}=3.87\,\mu_{B}$ for $S=3/2$, this moment is fully consistent with a $\mathrm{Co}^{2+}$ configuration once hybridization with the conduction band is taken into account, as demonstrated by DFT calculations [36, 23]. The underdoped compound Co0.22TaS2 exhibits an effective magnetic moment of $\sim 2.6\,\mu_{B}$ obtained from the Curie-Weiss fit, close to the spin-only value $2.83\,\mu_{B}$ for $S=1$, indicating a reduced effective spin state, consistent with a lower charge transfer. The lower charge transfer in underdoped compound implies that the lowest unoccupied Co orbital taken into account in our CPT is now below Fermi level, leaving only two remaining Co $d$ orbitals above the Fermi level. This influences the energy spectrum, making it more similar to the Ni intercalated case where $\beta$-feature was much weaker and shallower[23]. Furthermore, the disorder inevitably present in Co0.22TaS2, as evidenced by the negative temperature dependence of the electrical resistivity and relatively broad ferromagnetic onset, predominantly affects the Co-derived electronic states, suppressing the coherence required for the formation of a well-defined $\beta$-feature. The remaining bands are TaS2-derived and are much less affected by disorder, retaining sufficient coherence to remain observable in ARPES, although they appear significantly broadened. A difference is also seen in the shape of the central Fermi surface presented in Fig. 4(c), which in Co0.22TaS2 exhibits a hexagonal form, similar to that observed in Ni1/3NbS2 [23], in contrast to the circular Fermi surfaces of 33% Co intercalates [10, 11].
IV Conclusion
We have investigated the electronic structure and Fermi surface of Co1/3TaS2 and underdoped Co0.22TaS2 using ARPES combined with advanced theoretical modeling of electronic correlations. Both compounds exhibit strong similarities to the parent 2H-TaS2, however, in the 33% intercalated compound a shallow electron pocket, the so-called $\beta$-feature, emerges as a clear deviation from a rigid-band shift. Unlike previous interpretations that ascribed this feature to the surface effects, cluster perturbation theory (CPT) modeling reproduces the $\beta$-feature and demonstrates that it originates from strong electronic correlations and is strongly influenced by the intercalated cobalt orbitals.
These results build on our earlier indications in Co1/3NbS2, where tight-binding modeling of DFT results combined with a slave boson approach suggested a correlation driven origin of the $\beta$-feature. Independently, DMFT calculations reported in the literature [37] pointed in a similar direction. In the present study, by employing CPT, we provide the first quantitative modeling that reproduces the experimental ARPES spectra of Co1/3TaS2, thereby directly demonstrating that strong electronic correlations are essential for understanding the low energy electronic structure of intercalated transition-metal dichalcogenides.
Furthermore, ARPES measurements on underdoped Co0.22TaS2 reveal that the $\beta$-feature depends sensitively on electron doping, the orbital character of intercalant, and intrinsic disorder associated with the intercalated Co atoms. Our findings underscore the necessity of incorporating strong correlation effects in future theoretical studies of intercalated transition metal dichalcogenides.
Acknowledgements
We acknowledge insightful discussions with Eduard Tutiš. This work has been partly supported by the Croatian Science Foundation under the project numbers IP-2020-02-9666, UIP-2019-04-2154 and IP-2022-10-3382, by the project: Ground states in competition - strong correlations, frustration and disorder - FrustKor financed by the Croatian Government and the European Union through the National Recovery and Resilience Plan 2021-2026 (NPOO) and by the project Cryogenic Centre at the Institute of Physics - KaCIF co-financed by the Croatian Government and the European Union through the European Regional Development Fund-Competitiveness and Cohesion Operational Programme (Grant No. KK.01.1.1.02.0012). The research has been performed at the Elettra at the APE-LE beamline (proposal No. 20235254). NB acknowledges support by the project CeNIKS co-financed by the Croatian Government and the European Union through the European Regional Development Fund Competitiveness and Cohesion Operational Program (Grant No. KK.01.1.1.02.0013). The work at TU Wien was supported by the Austrian Science Fund (FWF) [10.55776/F86; 10.55776/P35945].
Appendix A. Wannier functions and Hamiltonian
Figure 5: Comparison of the DFT band structure (dotted lines) with the band structures obtained from the Wannier Hamiltonians, (solid lines). The upper panel shows the band tructure constructed using 52 Wannier functions per unit cell, while the lower panel displays the result obtained with a reduced basis of 7 Wannier functions per unit cell.
[…]
@@ Appendix A. Wannier functions and Hamiltonian
The idea of using Wannier functions is to develop a Hubbard like model in which the Hubbard interaction on the intercalated atoms will be treated more accurately compared to DFT calculations. The Wannier functions can be extracted from DFT by using software package Wannier90 [41]. We have done two versions of Wannier calculations. The first one, that is focused on the energy range from $E_{\rm F}$-7.7 eV to $E_{\rm F}$+1.5 eV, is based on 52 Wannier function within the unit cell of CoTa3S6. The space localized nature of Wannier functions makes possible to create a tight-binding like Hamiltonian which can accurately reproduce DFT band structure within given energy range. The result is shown in Fig. 5 in the upper panel.
@@
The second, simplified version of the Wannier calculation, is developed to describe the energy range just around the Fermi level, from $E_{\rm F}$-1.0 eV to $E_{\rm F}$+1.5 eV as this is the energy range where $\beta$-feature is observed. It is based on 7 Wannier functions in the unit cell. One may see that corresponding Hamiltonian can accurately reproduce 7 DFT bands around Fermi level, (the lower panel of Fig. 5). The top most Wannier band that is colored with red and is entirely above the Fermi level, has the largest contribution from lowest in energy empty Co $d$-state. The center of this Wannier function is located on Co atom. This Wannier function can be identified as an empty upper Hubbard state while the lower Hubbard state, that is occupied with an electron, is much lower in energy (around 6 eV) and it is outside the energy range in consideration. If this is a dispersion from spin up electrons, the center of Wannier function is located on the spin down Co ion, and vice versa. The centers of these Wannier functions depend on the electron spin in consideration. In spite of this differences between different spins, the band spectra for both spin projections are alike. The remaining 6 lower in energy bands have the largest contribution from Ta d-states. These Ta Wannier functions are not located on the Ta atoms, but in the points between three Ta atoms, in the space points not surrounded with sulphur atoms from above and below. These Wannier functions are shared with three neighboring Ta atoms. Their positions are independent on the spin projection. The most prominent terms of the simplified Wannier Hamiltonian are listed in Table.1. The largest hopping term in the Wannier Hamiltonian is between Co Wannier and Ta Wannier functions.

### v2

@@ III Results and discussions
The most prominent difference from the standard DFT+U results (for details see Fig. 6 shown in Appendix B) is the emergence of a shallow electron-like band with its minimum located at the $\overline{\mathrm{K}}$} point, clearly visible on the right side in Fig. 2(b). In the corresponding Fermi surface map (Fig. 2(c)), this band forms a triangular contour around the $\overline{\mathrm{K}}$} point, closely reproducing the experimentally observed $\beta$ feature. The orbital character analysis indicates that this state has a substantial Co contribution. Its appearance only within the CPT framework, which treats the Hubbard $U$ beyond the mean-field level, provides strong evidence that the $\beta$ feature originates from strong electronic correlations, in line with our earlier conjecture. In contrast, the remaining bands in the CPT calculation are very similar to those from DFT+U, which already provides a satisfactory description of the corresponding experimental spectra. An exception is the spectral weight around 0.8 eV binding energy, which originates from the highest S-derived band (see Appendix B) and is not captured by our reduced CPT model.
Figure 4: (Color online) (a) ARPES intensity plot of Co0.22TaS2 measured at $h\nu=72$ eV and 20 K. (b) Fermi surface along the $k_{z}$ direction, obtained by varying the photon energy from 57 to 90 eV in 3 eV steps at 20 K. (c) Fermi surface measured at $h\nu=63$ eV and 20 K.
Fig. 3 (a) shows the Fermi surface map of Co1/3TaS2 along the $k_{z}$ direction, obtained by varying the photon energy from 30 to 90 eV in small steps. The data were processed using the same procedure as described in Refs. [9, 22]. The $k_{z}$ dispersion plot reveals a two-dimensional (2D-like) Fermi surface, $\alpha_{2}$, visible at $k_{x}\approx\pm 0.3,\text{\AA }^{-1}$, which does not disperse with $k_{z}$ and shows only intensity variations. A faint additional 2D-like Fermi surface, $\alpha_{1}$, can also be observed near $k_{x}\approx\pm 0.15,\text{\AA }^{-1}$, although it is too weak to be traced across all $k_{z}$ values.
Both features arise from the Ta $d$ bands of 2H-TaS2: the outer band ($\alpha_{2}$) corresponds to the antibonding state, while the inner ($\alpha_{1}$), weaker one corresponds to the bonding state. DFT modeling predicts that the bonding band should exhibit noticeable $k_{z}$ dispersion and cross the Fermi level between the $\Gamma$ and A points. This explains why the $\alpha_{1}$ band is not observed across all $k_{z}$ values in Fig. 3(a). Indeed, the ARPES intensity plots measured with photon energies between $h\nu=38$-50 eV (for example see Fig. 6(a), presented in Appendix B) show the bonding band ($\alpha_{1}$) below the Fermi level. This photon energy range corresponds to $k_{z}$ values below 4 $\text{\AA }^{-1}$ in Fig. 3(a), which accounts for the disappearance of $\alpha_{1}$ at low $k_{z}$ values.
This $k_{z}$ dispersion agrees with the previously observed enhancement of $c$-axis electronic transport in TMDs upon Co intercalation, as discussed for Co1/3NbS2 in Refs. [9, 36]. The $\beta$ feature appears in Fig. 3(a) on both sides of the $\overline{\mathrm{K}}$ point, with its intensity increasing at higher photon energies, consistent with the behavior seen in Co1/3NbS2.
Fig. 3(b) shows the same Fermi surface as in panel (a), but calculated using the CPT method. The $\beta$ feature is again visible on both sides of the K point. Ta $d$ bands are also visible with antibonding band showing virtually no dispersion while bonding band exhibits strong $k_{z}$ dependence. We also notice the characteristic periodicity of two Brillouin zones resulting from the fact that there are two Co and two Ta layers per one unit cell. A similar effect was already reported by Weber et al. [52] in NbSe2.
To investigate the effect of reduced Co concentration on the electronic structure, ARPES measurements were performed on the underdoped Co0.22TaS2. The results are shown in Fig. 4. Notably, the band structure and Fermi surface closely resemble a rigid-band-shifted 2H-TaS2 structure, with the $\beta$ feature absent. By varying the incident photon energy from 57 to 90 eV, in Fig. 4(b) we mapped the Fermi surface along the $k_{z}$ direction. It exhibits a quasi-2D character with only a small modulation, and no $\beta$ feature is observed within the experimental photon energy range.
The central Fermi surface $\alpha$, around the $\mathrm{\Gamma}$ point, is observed at $k_{x}\sim 0.5~\text{\AA }^{-1}$, very close to that in 2H-TaS2 [32], indicating a much smaller band shift compared to the 33% intercalated compound and suggesting even lower charge transfer.
The same indication comes from the magnetic moment of Co1/3TaS2 compared to Co0.22TaS2. Co1/3TaS2 exhibits an effective magnetic moment of $\sim 3.3\,\mu_{B}$ similar to Co1/3NbS2 [36]. Despite being reduced compared to the spin-only value $\mu_{\mathrm{eff}}=3.87\,\mu_{B}$ for $S=3/2$, this moment is fully consistent with a $\mathrm{Co}^{2+}$ configuration once hybridization with the conduction band is taken into account, as demonstrated by DFT calculations [36, 22]. The underdoped compound Co0.22TaS2 exhibits an effective magnetic moment of $\sim 2.6\,\mu_{B}$ obtained from the Curie-Weiss fit, close to the spin-only value $2.83\,\mu_{B}$ for $S=1$, indicating a reduced effective spin state, consistent with a lower charge transfer. The lower charge transfer in underdoped compound implies that the lowest unoccupied Co orbital taken into account in our CPT is now below Fermi level, leaving only two remaining Co $d$ orbitals above the Fermi level. This influences the energy spectrum, making it more similar to the Ni intercalated case where $\beta$ feature was much weaker and shallower[22]. Furthermore, the disorder inevitably present in Co0.22TaS2, as evidenced by the negative temperature dependence of the electrical resistivity and relatively broad ferromagnetic onset, predominantly affects the Co-derived electronic states, suppressing the coherence required for the formation of a well-defined $\beta$ feature. The remaining bands are TaS2-derived and are much less affected by disorder, retaining sufficient coherence to remain observable in ARPES, although they appear significantly broadened. A difference is also seen in the shape of the central Fermi surface presented in Fig. 4(c), which in Co0.22TaS2 exhibits a hexagonal form, similar to that observed in Ni1/3NbS2 [22], in contrast to the circular Fermi surfaces of 33% Co intercalates [9, 10].
IV Conclusion
We have investigated the electronic structure and Fermi surface of Co1/3TaS2 using ARPES combined with advanced theoretical modeling of electronic correlations. In this 33% intercalated compound, a shallow electron pocket, the so-called $\beta$ feature, emerges as a clear deviation from a rigid-band shift, while it is notably absent in the underdoped Co0.22TaS2, highlighting the sensitivity of this feature to the intercalant concentration. To our knowledge, this represents the first experimental observation of the $\beta$ feature in an intercalated TaS2 system. Unlike previous interpretations in Co1/3NbS2, where this feature was ascribed to surface effects, our CPT modeling of Co1/3TaS2 reproduces the $\beta$ feature and demonstrates that it originates from strong electronic correlations, strongly influenced by the intercalated cobalt orbitals.
These results build on our earlier indications in Co1/3NbS2, where tight-binding modeling of DFT results combined with a slave-boson approach suggested a correlation-driven origin of the $\beta$ feature. Independently, DMFT calculations reported in the literature [37] pointed in a similar direction. In the present study, by employing CPT, we provide the first quantitative modeling that reproduces the experimental ARPES spectra of Co1/3TaS2, thereby directly demonstrating that strong electronic correlations are essential for understanding the low-energy electronic structure of intercalated transition-metal dichalcogenides.
Our study further demonstrates that the $\beta$ feature depends sensitively on the orbital character of the intercalant and the intrinsic disorder associated with the Co ions. These results underscore the necessity of incorporating strong correlation effects in future theoretical studies of intercalated transition-metal dichalcogenides.
Acknowledgements
We acknowledge insightful discussions with Eduard Tutiš. This work has been partly supported by the Croatian Science Foundation under the project numbers IP-2020-02-9666, UIP-2019-04-2154 and IP-2022-10-3382, by the project: Ground states in competition - strong correlations, frustration and disorder - FrustKor financed by the Croatian Government and the European Union through the National Recovery and Resilience Plan 2021-2026 (NPOO) and by the project Cryogenic Centre at the Institute of Physics - KaCIF co-financed by the Croatian Government and the European Union through the European Regional Development Fund-Competitiveness and Cohesion Operational Programme (Grant No. KK.01.1.1.02.0012). The research has been performed at the Elettra at the APE-LE beamline (proposal No. 20235254). NB acknowledges support by the project CeNIKS co-financed by the Croatian Government and the European Union through the European Regional Development Fund Competitiveness and Cohesion Operational Program (Grant No. KK.01.1.1.02.0013). The work at TU Wien was supported by the Austrian Science Fund (FWF) [10.55776/F86; 10.55776/P35945]. This work has been partly performed in the framework of the nanoscience foundry and fine analysis (NFFA-MUR Italy Progetti Internazionali) facility. The research leading to this result has been co-funded by the project NEPHEWS under Grant Agreement No 101131414 from the EU Framework Programme for Research and Innovation Horizon Europe
Appendix A. Wannier functions and Hamiltonian
Figure 5: (Color online) Comparison of the DFT band structure (dotted lines) with the band structures obtained from the Wannier Hamiltonians, (solid lines). The upper panel shows the band tructure constructed using 52 Wannier functions per unit cell, while the lower panel displays the result obtained with a reduced basis of 7 Wannier functions per unit cell.
@@ Appendix A. Wannier functions and Hamiltonian
The idea of using Wannier functions is to develop a Hubbard like model in which the Hubbard interaction on the intercalated atoms will be treated more accurately compared to DFT calculations. The Wannier functions can be extracted from DFT by using software package Wannier90 [41]. We have performed two versions of Wannier calculations. The first one, focused on the energy range from the Fermi level ($E_{\mathrm{F}}$) - 7.7 eV to $E_{\mathrm{F}}$ + 1.5 eV, is based on 52 Wannier functions within the unit cell of Co1/3TaS2. The space localized nature of Wannier functions makes possible to create a tight-binding like Hamiltonian which can accurately reproduce DFT band structure within given energy range. The result is shown in Fig. 5 in the upper panel.
@@
The second, simplified version of the Wannier calculation, is developed to describe the energy range just around the Fermi level, from $E_{\rm F}$-1.0 eV to $E_{\rm F}$+1.5 eV as this is the energy range where $\beta$ feature is observed. It is based on 7 Wannier functions in the unit cell. One may see that corresponding Hamiltonian can accurately reproduce 7 DFT bands around Fermi level, (the lower panel of Fig. 5). The top most Wannier band that is colored with red and is entirely above the Fermi level, has the largest contribution from lowest in energy empty Co $d$-state. The center of this Wannier function is located on Co atom. This Wannier function can be identified as an empty upper Hubbard state while the lower Hubbard state, that is occupied with an electron, is much lower in energy (around 6 eV) and it is outside the energy range in consideration. If this is a dispersion from spin up electrons, the center of Wannier function is located on the spin down Co ion, and vice versa. The centers of these Wannier functions depend on the electron spin in consideration. In spite of this differences between different spins, the band spectra for both spin projections are alike. The remaining 6 lower in energy bands have the largest contribution from Ta $d$-states. These Ta Wannier functions are not located on the Ta atoms, but in the points between three Ta atoms, in the space points not surrounded with sulphur atoms from above and below. These Wannier functions are shared with three neighboring Ta atoms. Their positions are independent on the spin projection. The most prominent terms of the simplified Wannier Hamiltonian are listed in Table.1. The largest hopping term in the Wannier Hamiltonian is between Co Wannier and Ta Wannier functions.